OpenStudy (anonymous):

if i raise 4 to the 214th power, whats the units digits?of the result?

OpenStudy (anonymous):

You'll want to use modular arithmetic here. In particular, you want to know: $$4^{214}\pmod{10}$$Is this something you have seen before?

OpenStudy (anonymous):

If not thats fine, this can be explained without using modular arithmetic. Its just a little easier to explain with it.

OpenStudy (anonymous):

i plugged 4^214 the answer is 6.931.674e the unite digit should be 6

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

maybe im mistaking something

OpenStudy (anonymous):

Just try the first few powers and see if you spot a pattern in the ones digits: $$4^1=4$$$$4^2=16$$$$4^3=64$$$$4^4=256$$$$4^5=1024$$$$4^6=4096$$

ganeshie8 (ganeshie8):

\[\large 4^{214} = 16^{107}\] 16 maybe easier to work

OpenStudy (anonymous):

thanks